### v2

# Evaluating LLM-based Approaches to Legal Citation Prediction: Domain-specific Pre-training, Fine-tuning, or RAG? A Benchmark and an Australian Law Case Study

Abstract
Large Language Models (LLMs) have demonstrated strong potential across legal tasks, yet the problem of legal citation prediction remains under-explored. At its core, this task demands fine-grained contextual understanding and precise identification of relevant legislation or precedent. We introduce the AusLaw Citation Benchmark, a real-world dataset comprising $55k$ Australian legal instances and $18,677$ unique citations which to the best of our knowledge is the first of its scale and scope. We then conduct a systematic benchmarking across a range of solutions: (i) standard prompting of both general and law-specialised LLMs, (ii) retrieval-only pipelines with both generic and domain-specific embeddings, (iii) supervised fine-tuning, and (iv) several hybrid strategies that combine LLMs with retrieval augmentation through query expansion, voting ensembles, or re-ranking. Results show that neither general nor law-specific LLMs suffice as stand-alone solutions, with performance near zero. Instruction tuning (of even a generic open-source LLM) on task-specific dataset is among the best performing solutions. We highlight that database granularity along with the type of embeddings play a critical role in retrieval-based approaches, with hybrid methods which utilise a trained re-ranker delivering the best results. Despite this, a performance gap of nearly $50\%$ remains, underscoring the value of this challenging benchmark as a rigorous test-bed for future research in legal-domain. [^1] The data, code, and trained LLMs and re-rankers are available at: https://auslawbench.github.io
Evaluating LLM-based Approaches to Legal Citation Prediction: Domain-specific Pre-training, Fine-tuning, or RAG?
A Benchmark and an Australian Law Case Study
Jiuzhou Han     Paul Burgess     Ehsan Shareghi  Department of Data Science & AI, Monash University  Faculty of Law, Monash University first.last@monash.edu
1 Introduction
Recent advancements in utilising Large Language Models (LLMs) for legal domain have shown promising results across various tasks. For instance, Pont et al. (2023) leveraged LLMs for generating summaries of judicial decisions, identifying legal issues, decision-making criteria, and specifying keywords. Deroy et al. (2024) revealed that LLMs outperform extractive summarisation methods in quality metrics but suffer from inconsistencies and hallucinations, highlighting the importance of human-in-the-loop approaches for improved reliability. Jiang and Yang (2023) underscored that fine-tuning LLMs demonstrates state-of-the-art performance in legal judgment prediction, while Peng and Chen (2024) showed retrieval-augmentation leads to improved accuracy by integrating external knowledge, particularly for complex charges. Hou et al. (2024a) proposed a model to detect deviations between an AI-generated legal analysis and human as a way of quantifying their reliabilities.
While these advancements highlight LLMs’ transformative potential in legal applications, challenges like ensuring factual accuracy, handling diverse tasks, and mitigating inherent issues such as hallucination remain. For instance, Dahl et al. (2024) reports that even state-of-the-art LLMs hallucinate between 69-88% of responses to legal queries, while Magesh et al. (2024) highlights that hallucination issue is mitigated with specialisation of tools but still persists as an unresolved issue.
In this paper, we report progress with respect to a less-explored task in the legal domain, Legal Citation Prediction. Citations in legal cases, as in academic writing, serve two purposes: first, information necessary to locate, read, and verify the material; and second, information about the authority of the source is conveyed (Axel-Lute, 1982). The second is particularly crucial in legal cases in common law jurisdictions, as in those systems most decision makers are required to follow previous decisions (Schauer, 1987). Accordingly, whilst precedent plays a fundamental role in determining how courts behave and, therefore, in how societies function, citations to, of, and between authorities are the way in which precedent is communicated. In determining matters in court, judges use citations to give weight and authority to their decisions and also, crucially, to demonstrate that they are acting appropriately and are adhering to the legal precedent that already exists. When doing so, a judge or a panel of judges will often state a legal rule – as a principal or a proposition – and then support the existence of that rule with a citation to another source; frequently a prior court decision. In other words, judges making determinations today rely on citations to not only determine legal questions in court but also to show that they are acting within their (the judges’) lawful role.
Formally, the citation prediction task can be defined as follows: Given a passage, the goal is to identify the correct legislation or precedent that applies and needs to be cited (i.e., to predict or retrieve the correct [CITATION]). The following examples illustrate the citation prediction tasks considered in this paper:
Example 1.
Query: The distinction between a genuine offer of compromise and a demand to capitulate has to be recognised. See the discussion in [CITATION].
Answer: Leichhardt Municipal Council v Green [2004] NSWCA 341
Example 2.
Query: The Tribunal is satisfied that the applicant does not fulfil the requirements of section 139(a) of the National Law, in that she lacks the mental capacity to practise medicine, as was considered in [CITATION].
Answer: Lindsay v Health Care Complaints Commission [2010] NSWCA 194
Example 3.
Query: Whilst it is suggested that the offender’s mother and grandmother have difficulty paying rent without the offender’s assistance, there is no evidence of how they support themselves or their financial circumstances. There is no evidence of hardship that might meet the ‘truly, wholly or highly exceptional’ standard referred to in [CITATION].
Answer: Jinnette v R [2012] NSWCCA 217
We aim to compare, develop, and explore different solutions for the citation prediction task. We release, AusLaw Citation Benchmark, a real dataset of $55k$ instances specific to Australian law, covering $18,677$ unique citations. To the best of our knowledge this is the first dataset of this scale and scope. We then conduct a thorough comparison along the following comprehensive dimensions:
  • Prompting general purpose LLMs (i.e., GPT-4o (OpenAI, 2024; Achiam et al., 2023), Claude Sonnet 3.5 (Anthropic, 2024), LLaMA-3.1-70B-instruct (Grattafiori et al., 2024), Command R+ (Cohere, 2024))
  • Prompting law-specialised pre-trained LLMs (i.e., SaulLM-7B-instruct (Colombo et al., 2024b), SaulLM-54B-instruct (Colombo et al., 2024a))
  • Retrieval-only setup with vectorised DB using general-purpose (i.e., text-embedding-3-large) and law-specific embeddings (i.e., AusLaw-embedding-v1.0)
  • Instruction fine-tuning LLMs (i.e., SaulLM-7B, and LLaMA-3.1-8B) for the citation prediction task
  • Different hybrid tactics that combine LLMs and retrieval systems (i.e., retrieval-augmented generation, query expansion, voting ensemble, and specialised re-rankers)
While we aim to push all these solutions to their limits to better understand their effectiveness and limitations, a near $50\%$ gap is remained to be filled. We hope this benchmark and the baselines and methods investigated in this paper to encourage future developments in this important and challenging task.
2 Legal NLP in the Era of LLMs
While a satisfying review demands a separate work, we attempt to provide a brief overview of existing work on the intersection of Law and LLMs. We also acknowledge that this only provides an overview of a small subset of the broader space of research in Legal NLP.
The intersection of NLP and the legal domain has given rise to a wide array of research tackling tasks such as legal text classification, retrieval, summarisation, question answering, and reasoning. As legal texts are often complex and domain-specific, specialised benchmarks and models have become essential to evaluate and improve NLP performance in this high-stakes field.
2.1 Legal Benchmarks
Recent efforts in benchmarking legal NLP have produced a diverse ecosystem of datasets targeting legal reasoning, retrieval, and comprehension. Benchmarks like LEGALBENCH Guha et al. (2024), LawBench Fei et al. (2023), and LexEval offer broad, multi-task evaluations across cognitive levels and languages, revealing persistent limitations in current LLMs’ legal reasoning abilities. Task-specific benchmarks like BLT Blair-Stanek et al. (2023) and MAUD Wang et al. (2023) assess fundamental skills like basic legal text navigation and merger agreement comprehension, while CUAD Hendrycks et al. (2021) and the statutory reasoning Holzenberger et al. (2020) in tax law dataset bring attention to contract review and rule-based reasoning tasks demanding deep understanding of legal semantics and structure. LexGLUE Chalkidis et al. (2022) and LEGAL-BERT Chalkidis et al. (2020) contribute to general legal language understanding and adaptation of pretrained models for legal tasks, while LegalBench-RAG Pipitone and Alami (2024) and CLERC Hou et al. (2024b) evaluate retrieval and generation in legal writing tasks. Newer benchmarks such as LegalAgentBench Li et al. (2024a) assess autonomous legal agents through multi-step tasks and tool use in real-world scenarios, while LegalHalBench Hu et al. (2025) introduces fine-grained metrics and datasets to evaluate and reduce hallucination in legal QA systems. Region-specific datasets like IL-TUR Joshi et al. (2024) for Indian law and LexEval Li et al. (2024b) for Chinese legal systems underscore the need for jurisdictionally relevant evaluations. We propose a new benchmark focused on legal citation prediction.
2.2 Law-Specialised Large Language Models
The development of law-specialised LLMs has gained momentum in response to the distinctive demands of legal language, which is domain-specific, syntactically complex, and semantically nuanced. Early adaptation efforts such as LEGAL-BERT Chalkidis et al. (2020) explored various strategies for pretraining BERT models on legal corpora, revealing that full domain-specific pretraining outperforms generic approaches. More recent models like Lawma Dominguez-Olmedo et al. (2024) and LawLLM Shu et al. (2024) demonstrate the effectiveness of fine-tuned and multi-task architectures in addressing legal classification, retrieval, and judgment prediction tasks within the United States legal system. The SaulLM Colombo et al. (2024b, a) family is the first (and only) open-source Law-specialised LLMs ranging from 7B to 141B parameters, using extensive legal corpora (500B tokens), instruction tuning, and preference alignment to achieve state-of-the-art results across multiple legal benchmarks. These models highlight the benefits of legal specialisation.
Nevertheless, the presence of domain-specialised LLMs only reduce the semantic gap required for the domain and still require (as we will show later in the experiments) task-specific training to be competent for fine-grained tasks such as legal citation prediction. Additionally, we argue (and empirically show) that pretraining on a mixture of data from various jurisdictions and countries does not provide a reliable understanding of the legal context specific to each of those jurisdictions. To bridge this gap, we develop an Australia-specific legal LLM focused on tasks such as legal citation prediction—thereby supporting jurisdiction-aware applications and advancing the broader field of legal artificial intelligence.
3 AusLaw Citation Benchmark
To build our benchmark, we use the NSW Caselaw section of the Open Australian Legal Corpus. [^2] https://huggingface.co/datasets/umarbutler/open-australian-legal-corpus We identified $82,530$ citations that specifically referred to a case within the dataset. For each citation ($c_{k}$), we extracted the sentence in which the citation appeared (denoted as $S^{i}_{c_{k}}$), along with the preceding sentence (denoted as $S^{i-1}_{c_{k}}$). We further utilised an LLM to generate an auxiliary description of $c_{k}$, based on (${{Full\ Text}_{c_{k}}},S^{i}_{c_{k}},S^{i-1}_{c_{k}}$) where ${Full\ Text}_{c_{k}}$ refers to the full text of the cited case. We manually checked a subset of the LLM-generated descriptions for quality assurance and optimizing the prompt wording. An example of a generated RoC is provided below:
$S^{i-1}_{c_{k}},S^{i}_{c_{k}}$: In considering what if any orders should be made in regards to the surface roots, I am not satisfied to the level required by s 10(2) of the Act, that there is any real likelihood of injury arising from those roots. Craig J in Leichhardt Municipal Council v Green [2004] NSWCA 341, considered that ’something more than a theoretical possibility is required in order to engage the power under the Trees Act’.
CITATION $c_{k}$: Leichhardt Municipal Council v Green [2004] NSWCA 341
Generated Reason-of-Citation (RoC): The cited case is referenced to establish the standard required to demonstrate a likelihood of injury under the Trees Act.
We refer to this description as Reason-of-Citation (RoC). [^3] In this process, we tasked the LLM to not generate any RoC (i.e., to generate NOT ENOUGH INFORMATION) if the information was not adequate in $(S^{i}_{c_{k}},S^{i-1}_{c_{k}})$. These refused cases were not included in the data. See Table A.4 of Appendix A.4 for details on when this was triggered. Each reference to a citation $c_{k}$ in the data results in a unique new $\text{RoC}_{c_{k}}$. For each $c_{k}$, we denote its $M$ references as $\mathrm{RoC}_{c_{k}}^{1},\mathrm{RoC}_{c_{k}}^{2},\ldots,\mathrm{RoC}_{c_{k}}^{M}$. We will discuss later how the RoCs are used. This resulted in the final dataset of $55,005$ instances, covering $18,677$ unique citations. Within this set, 5% of the citations have been referenced at least 9 times, while 54% were cited only once. From this final set, we extracted $1k$ citations as test set, and used the rest for training (i.e., our instruction-tuned models). See Appendix A.2 for more details.
4 Methods
We investigate various methods under Open World and Closed World settings. The Open World setting places no restriction over the possible predictions from the system (i.e., similar to how an LLM functions in real-world), whereas the Closed World setting confines the output space to be from the set of $18,677$ citations present in the database (i.e., similar to a standard retrieval setup).
4.1 LLM-Only
We explored both existing LLMs (General purpose and Law-specialised) as well as our instruction-tuned LLMs.
Existing General Purpose LLMs.
For the general purpose LLMs we used GPT-4o (OpenAI, 2024; Achiam et al., 2023), Claude Sonnet 3.5 (Anthropic, 2024), LLaMA-3.1-70B-instruct (Grattafiori et al., 2024), and Command R+ (Cohere, 2024). When directly prompted with the query, all these LLMs demonstrated near-zero performance. To gain deeper insight, we leaked the RoCs along with the query text. This approach, while bypassing the task’s requirement of predicting the citation without access to any information about the case-to-be-cited, allowed us to estimate the upper-bound performance of these models. More on this will be discussed in Section 5.
Existing Law-specialised LLMs.
For law-specialised LLMs, we used SaulLM-7B-instruct (Colombo et al., 2024b), SaulLM-54B-instruct (Colombo et al., 2024a) which to the best of our knowledge are the only publicly available law LLMs for English to this date. [^4] We were unable to run SaulLM-141B-Instruct due to its substantial hardware requirements. For prompting, we followed the exact setting of general purpose LLM experiments (see Table A.4 of Appendix A.4).
Our Instruction Fine-tuned LLMs.
We instruction-tuned LLaMA-3.1-8B and SaulLM-7B-Base models on the training data. These resulting models are referred to as Cite-LLaMA-3.1-8B and Cite-SaulLM-7B in our experiments. At inference time, given the query ($S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}})$) the model first produces the $\texttt{RoC}_{c_{k}}$ and then $c_{k}$ (i.e., $p(c_{k},\texttt{RoC}_{c_{k}}\mid I,S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}});\theta)$:
$p(c_{k}\mid I,S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}}),\texttt{RoC}_{c_{k}};\theta)$ (1)
×$\quad\times p(\texttt{RoC}_{c_{k}}\mid I,S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}});\theta).$
where $\theta$ denotes LLM parameters, $I$ denotes the instruction, $\texttt{Mask}(S^{i}_{c_{k}})$ denotes $S^{i}_{c_{k}}$ with the citation to $c_{k}$ being masked. See Appendix A.3 for further details on training parameters and instruction detail and format (same instruction was used at inference).
Unlike the previous setups, for all experiments with the instruction-tuned models only the text of the query (i.e., no RoC) was provided and the model is instructed to predict both the $RoC_{c_{k}}$ and the citation $c_{k}$. This presents a challenging setup as the LLM needs to predict a citation in the Open World setting solely by its parametric knowledge and the brief information provided in the query text.
4.2 Retrieval-Only
We investigated retrieval along two axes: embeddings, and granularity of data to be indexed in the vectorised database. The basic principle is a retrieval task where a query is matched against all records of a database, and the Top-k closest records are returned. The closeness is measured in the semantic space through embeddings. In all setups involving retrieval, unless stated otherwise, the search returns the Top-5 most relevant citations.
Embeddings.
For embeddings in the retrieval system (both for representation of queries and the vectorized database) the 3072 dimensional text-embedding-3-large [^5] https://openai.com/index/new-embedding-models-and-api-updates from OpenAI as generic embeddings, and the 384 dimensional vectors of AusLaw-embedding-v1.0  [^6] https://huggingface.co/adlumal/auslaw-embed-v1.0 as Australian law-specialised embeddings were used. [^7] While we used the largest dimensionality of embeddings available to measure the performance upper-bounds, there are directions of exploration which we did not pursue in this work: the impact of using various dimensionalities on the retrieval system, or alternative means of building dense or conventional types of indexes (i.e., Lucene/BM25).
Database Granularities.
To construct the database, three possibilities for representing each citation ($c_{k}$) were considered: (1) Full Case Text ($\text{Full Case}_{c_{k}}$), (2) Catchwords ($\text{Catchwords}_{c_{k}}$), and (3) Aggregation of all its corresponding $M$ RoCs (RoC Aggregations=concat($\text{RoC Aggregations}=\operatorname{concat}(\mathrm{RoC}^{1}_{c_{k}},\mathrm{RoC}^{2}_{c_{k}},\ldots,\mathrm{RoC}^{M}_{c_{k}})$).
We refer to these different settings as Index Granularity later in the experiments (i.e., Table 1). For each granularity level, we created a distinct database version for each embedding backbone. [^8] We use https://github.com/chroma-core/chroma for building the DB, and use the built-in cosine similarity for measuring closeness.
4.3 Hybrid of LLM and Retrieval
What follows can be regarded as various instances of Retrieval Augmented Generation (RAG). However, in our manuscript, we strictly refer to RAG as the scenario where the LLM generation is guided by retrieval, rather than retrieval being influenced by the generation.
Query Expansion.
[…]
Voting Ensemble
First, following the LLM-only setup, instruction-tuned LLMs (e.g., Cite-LLaMA-3.1-8B) generate the RoC’ and the citation. Next, the query is expanded into Text+RoC’ and fed into the retrieval system. If the LLM-generated citation appears in the Top-5 retrieval results, it is returned; otherwise, the Top-1 result from the retrieval system is returned. This is to leverage the benefit of LLM-only setup by reducing its hallucination (i.e., producing citations outside the corpus).
Retrieval Augmented Generation (RAG).
This follows a standard Retrieval-Augmented Generation (RAG) setup, where queries are sent to the database, and the Top-5 results are retrieved for the LLM to re-rank and select the Top-1 match. See Table A.4 of Appendix A.4 for prompt details.
Re-ranker.
We have four settings to train the re-ranker. The same data used for instruction fine-tuning was used for training the re-rankers. The input of a re-ranker is the text containing the missing citation and retrieved top-5 candidate citations and their corresponding RoC. The output is the gold citation. Since the RoC of each case has different length, to train a re-ranker, we needed to eliminate the effect of discrepancy in the length.
  • Setting 1 uses the first reason as the whole citation reason.
  • Setting 2 tasks GPT-4o-mini to merge the citation reasons of each case into a similar length text.
  • Setting 3 leverages the same merged citation reason as Setting 2. In addition, given 5 cases and their case texts, the GPT-4o-mini is tasked to generate the rational of why the gold one is cited among all 5 citations.
  • Setting 4 is based on retrieval. We retrieve the most similar citation reason from the citation reasons list as the citation reason of that case. Within this setting, there are two variations: 1) we use the retrieved top-1 citation reason as the output for the training; 2) we use the gold citation reason as the output for the training. We name these two variations as 4v1 and 4v2.
See Table 10 of Appendix A.3 for input and output details.
5 Experiments
                   LLM-only Approach: Direct zero-shot prompting
Type LLMs Query Output ACC@1 ACC@5
Open World General Purpose GPT-4o Text+RoC Top-5 Citations 0.1 0.1
Claude Sonnet 3.5 Text+RoC Top-5 Citations 15.5 16.8
Command R+ Text+RoC Top-5 Citations 0.0 0.0
LLaMA 3.1 70B Instruct Text+RoC Top-5 Citations 1.6 2.1
Law-specialised SaulLM-7B-Instruct Text+RoC Top-5 Citations 0.0 0.0
SaulLM-54B-Instruct Text+RoC Top-5 Citations 2.0 2.7
Citation-tuned (ours) Cite-SaulLM-7B Text RoC+Top-1 Citation 51.7 -
Cite-LLaMA-3.1-8B Text RoC+Top-1 Citation 46.2 -
         Retrieval-only Approach: Uses vectorised database and vectorised Query to retrieve Top-5
Embeddings Index Granularity Query Output ACC@1 ACC@5
text-embedding-3-large Full Cases Text Top-5 Citations 14.9 32
Catchwords Text Top-5 Citations 14.7 32.5
RoC Aggregations Text Top-5 Citations 27.1 53.8
AusLaw-embedding Full Cases Text Top-5 Citations 8.7 20.7
Catchwords Text Top-5 Citations 10.5 22.4
RoC Aggregations Text Top-5 Citations 29.5 54.5
[…]
Results are formatted as GPT-4o/SaulLM-54B-Instruct/Cite-LLaMA-3.1-8B/Cite-SaulLM-7B
Embeddings Index Granularity Query Output ACC@1 ACC@5
text-embedding-3-large Full Cases Text Top-5 Citations 14.3/14.4/17.1/17.4 31.1/31.4/34.3/34.4
Catchwords Text Top-5 Citations 15.3/15.5/15.0/15.8 33.1/33.1/33.4/33.9
RoC Aggregations Text Top-5 Citations 29.6/28.6/34.9/35.1 56.7/56.1/60.0/60.4
AusLaw-embedding Full Cases Text Top-5 Citations 9.0/9.5/11.7/12.4 21.1/21.3/24.2/26.0
Catchwords Text Top-5 Citations 10.2/10.9/11.0/11.4 23.5/24.6/24.3/24.4
RoC Aggregations Text Top-5 Citations 32.2/30.4/33.5/34.7 55.8/54.2/55.6/56.5
         (Hybrid Approach) Voting Ensemble: Returns LLM’s citation if in the Top-5 of retrieval; otherwise, returns the retrieval’s Top-1
Results are formatted as Cite-LLaMA-3.1-8B/Cite-SaulLM-7B
Embeddings Index Granularity Query Output ACC@1 ACC@5
text-embedding-3-large RoC Aggregations Text Top-5 Citations 47.3/48.2 -
AusLaw-embedding RoC Aggregations Text Top-5 Citations 43.6/45.3 -
         (Hybrid Approach) RAG: Given the Query, first retrieves Top-5 citations, and uses GPT-4o to pick the best
Embeddings Index Granularity Query Output ACC@1 ACC@5
text-embedding-3-large Full Cases Text Top-5 Citations 16.5 -
Catchwords Text Top-5 Citations 21.7 -
RoC Aggregations Text Top-5 Citations 42.2 -
AusLaw-embedding Full Cases Text Top-5 Citations 10.2 -
Catchwords Text Top-5 Citations 17.1 -
RoC Aggregations Text Top-5 Citations 42.9 -
         (Hybrid Approach) Re-ranker: Given the Query, first retrieves Top-5 citations, and use trained re-ranker to pick the best
Results are formatted as setting1/2/3/4v1/4v2
Embeddings Index Granularity Query Output ACC@1 ACC@5
text-embedding-3-large RoC Aggregations Text Top-5 Citations 20.2/26.9/21.7/50.4/50.9 -
AusLaw-embedding RoC Aggregations Text Top-5 Citations 20.6/28.5/22.9/50.2/51.2 -
[…]
Results are formatted as setting1/2/3/4v1/4v2
Embeddings Index Granularity Query Output ACC@1 ACC@5
text-embedding-3-large RoC Aggregations Text Top-5 Citations 21.4/28.5/23.8/51.7/$52.1^{*}$
AusLaw-embedding RoC Aggregations Text Top-5 Citations 21.2/28.3/23.1/50.8/51.4 -
Table 1: Citation accuracy under various methods, including ours. For results marked by $\ddagger$, we used gpt-4o-mini due to cost. Best result is marked with . The rows are colour-coded for readability and easier comparison of settings referenced in the discussion §5.1. Text denotes ($S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}})$), see §4.1 for notations.
LoRA (Hu et al., 2022) and 8-bit quantization was used for the instruction tuning stage. For details regarding the instruction fine-tuning step please refer to Table 3 of Appendix A.3. As evaluation metrics, Accuracy@1 and Accuracy@5 are used.
5.1 Results
The main results are presented in Table 1. We structure our discussion of results along the comparisons outlined in the introduction (§1).
Pre-training on general text vs. law-specialised text.
The results of the LLM-only experiments (1st panel of Table 1) reveal that pre-training alone is insufficient for achieving satisfactory performance in the citation prediction task. Neither general-domain nor law-specialised models demonstrated reasonable accuracy. For example, Claude Sonnet 3.5 achieved only 15% accuracy in predicting citations when provided with both the query text and the RoC. In contrast, both the 7B and 54B variants of SaulLM performed even worse, achieving 2% accuracy or less, despite being explicitly pre-trained on the Open Australian Legal Corpus (including the NSW Caselaw subset used in this study) and other law-specific datasets. This highlights that mere domain-specific pre-training of relatively large (i.e., SaulLM-54B) is not sufficient for fine-grained tasks.
Also it is noteworthy that out of the 94-520B tokens used in pre-training of such Lawps-specialised LLMs, only 0.5B tokens were covering Australian jurisdiction (see Table 1s of Colombo et al. (2024b) and Colombo et al. (2024a)). This opens a natural question: whether specific pre-training solely on that segment of 0.5B tokens will give a more jurisdictional chance to the LLM in our task. While a proper investigation of this will require a substantial GPU support, we investigate this question for the 7B-8B LLM scales in Subsection 5.2.
Pre-training vs. Instruction tuning.
A significant performance boost is observed when comparing pre-trained LLMs to their instruction-tuned counterparts on our training data. For example, the SaulLM-7B model’s performance jumps from 0% to 51.7% in the more challenging Text-only query setup. Similarly, LLaMA-3.1-8B, a general-purpose model, achieves 46.2% accuracy after instruction fine-tuning, surpassing its much larger pre-trained 70B counterpart. Notably, between the LLaMA and SaulLM backbones, instruction tuning on the same dataset and for the same number of epochs proves more effective for the domain-specific SaulLM. This underscores the critical importance of not only domain-specialised pre-training but also targeted fine-tuning on task-specific data and requirements.
As anticipated, Figure 1 demonstrates a negative correlation between predictive accuracy and the citation frequency in the dataset. Cases cited more than 100 times achieve 100% accuracy, while accuracy drops below 40% for cases cited 20 times or fewer. This reflects the challenge fine-tuning faces in accurately predicting less frequently cited cases (i.e., freq<20), where alternative approaches which integrate retrieval (i.e., retrieval with re-ranking, explored next) might offer a more reliable solution.
Figure 1: Average Accuracy and Number of Unique Cases for various citation frequency buckets. Accuracies per bin are based on the following settings from Table 1: Cite-SaulLM-7B (ACC@1: 51.7), RAG (ACC@1: 42.9), and Re-ranker (ACC@1: 52.1), and from Table 2: Cite-AusLawLLM-7B (ACC@1: 52.0).
Generic vs. domain-specialised embeddings.
When comparing the best results of each block that involves retrieval: red rows (corresponding to the generic text-embedding from OpenAI) vs. orange rows (corresponding to embeddings trained exclusively on the Open Australian Legal Corpus) in each block of results in Table 1, it is important to to acknowledge that this comparison is not entirely fair due to differences in embedding dimensionality (3072 vs. 384), the volume of data used, and the training algorithms employed. Nonetheless, we observe an overall intuitive pattern in favour of the domain-specialised AusLaw-embeddings, or at least being on-par with OpenAIs emeddings (while being 8×8$\times$8 × smaller in dimensionality). This highlights the benefits of domain-specialised embeddings and suggests a few promising future directions to explore: training larger embeddings tailored to the Australian legal domain, while we do not investigate this further, a trivial direction of exploration could be to build higher dimensional representations, as well as improving the geometric utilisations of the representation space (e.g., through isotropy Liu et al. (2021)). We leave further explorations of these to future work.
Database granularity matters.
Probing the results along the Index Granularity dimension indicates the significant role the granularity of database (index) plays in the accuracy of the retrieval system. Contrary to our expectations, Catchwords proved to be the least effective granularity, performing worse than Full Cases. By a substantial margin, the best performance was achieved with the RoC Aggregations granularity. This pattern is consistent for both the generic and law-specialised embeddings, and across all setups except for the Retrieval-only were Full Case leads to slightly better results compared with Catchwords.
Query expansion vs. voting ensemble vs. RAG.
Comparing all retrieval-based methods, excluding re-ranking, hybrid methods are consistently better than Retrieval-only. The Voting Ensemble is the best, followed by RAG and then Query Expansion. The superiority of Voting Ensemble highlights the advantages of combining the predictive quality of instruction-tuned LLMs with the robustness of a retrieval system. In contrast, RAG relies on query augmentation to guide the LLM’s predictions within the context of the Top-5 retrieved citations, while the Query Expansion method instead focuses on re-adjusting the query’s semantic space before searching the retrieval space.
What we can learn from ACC@5.
The general pattern across all experiments suggests the potential that lies within Top-5 retrieved citations. For instance, while the Query expansion results are not competitive at ACC@1, the promising 60.4% for Cite-SaulLM-7B for ACC@5 (the second row highlighted in red) suggests that with improved re-ranking of the Top-5 hits, there is potential to boost accuracy by up to 10%. We explore this next.
Re-ranker: the Gap of ACC@1 and ACC@5.
The ACC@5 results indicate that the correct citation is often present within the top five retrieved candidates. Building on this insight, the re-ranker’s goal is to enhance the likelihood of selecting the correct citation from these Top-5 candidates. The bottom two blocks in Table 1 demonstrate the performance of various re-ranking methods. Notably, the best-performing re-ranker boosts ACC@1 from 35.1% to 52.1%, significantly narrowing the gap toward the upper bound defined by ACC@5 (60.4%).
5.2 Australian Law Pre-training
SaulLM-7B was pre-trained on a 94B-token general-domain legal corpus comprising various legal sources, including 0.5B tokens from the Open Australian Legal Corpus. To investigate whether pre-training solely on Australian legal data is sufficient for building an effective Australia-specific legal LLM, we pre-train the same underlying vanilla Mistral-7B model and a LLaMA-3.1-8B model (used in SaulLM), on the 0.5B tokens of Australian Legal corpus. Following pre-training, we apply the same citation instruction-tuning procedure on the citation prediction task. While we only pre-train the underlying LLM for 5 epochs on the 0.5B tokens, due to infrastructure limitations, we observe promising benefits of jurisdictional pre-training. The results, presented in Table 2, show that Cite-AusLawLLM-7B slightly outperforms Cite-SaulLM-7B (both built on Mistral-7B), despite the latter being trained on a much larger 94B-token corpus. Furthermore, we observe that pre-training significantly boosts performance of the LLaMA-3.1-8B model after Australian Law pre-training pre-training (Cite-AusLawLLM-8B).
Model Law Data Accuracy
Cite-SaulLM-7B (Mixed Law Pre-training) 94B 51.7
Cite-AusLawLLM-7B (AusLaw Pre-training) 0.5B 52.0
Cite-LLaMA-3.1-8B (No Law Pre-training) - 46.2
Cite-AusLawLLM-8B (AusLaw Pre-training) 0.5B 50.0
Table 2: The results of different pre-training data (Mixed vs. AusLaw-specific) before doing instruction-tuning on citation prediction task. Data size is reported in tokens.
6 Conclusion
In this paper, we proposed a new benchmark, AusLaw Citation Benchmark, for legal citation prediction and examined various approaches of leveraging Large Language Models and retrieval systems, evaluating their effectiveness both independently and in combination. Our findings demonstrate that while pre-training large language models on general or even domain-specialised legal texts is a necessary starting point, it is far from sufficient for achieving satisfactory citation prediction accuracy in the Australian legal domain. The most contributing factor for improving performance lies in targeted instruction tuning with task-specific data, which dramatically boosts accuracy. Our experiments further reveal the importance of choosing the right embedding model and database granularity for retrieval, with results showing up to 70% variation in performance under different granularities. Among the retrieval-augmented methods, ensemble voting strategy stands out as the most effective, outperforming methods like RAG and query expansion. Furthermore, we show that training re-rankers effectively harnesses the untapped gains available in top-k retrieval accuracy.
While the above approaches offer varying degree of effectiveness, there remains a noticeable $50\%$ gap in performance that calls for further developments. We hope the proposed benchmark and the investigated approaches provide a reliable framework for development of more advanced solutions that intersect LLMs, fine-tuning, pre-training, and retrieval mechanisms.
Limitations
First, due to computational constraints, we limit most of our open-source experiments to 7B-parameter models and do not evaluate the performance of larger-scale models (i.e., SaulLM 141B), which may offer additional gains. Second, our citation prediction task is currently framed as a single-citation prediction problem, where each input maps to only one citation. A more realistic and challenging setting would involve predicting multiple citations per instance under a larger context. Additionally, in this paper we did not test the methods in the out-of-distribution setting which could further highlight other potential challenges in legal domain.
Appendix A Appendix
A.1 Three Examples of Catchwords
Case 1: CRIME – Appeals – Appeal against conviction – Unreasonable verdict – two counts of sexual offences – where applicant found guilty on one count and acquitted on the other – whether on all of the evidence it was open to the jury to be satisfied of the applicant’s guilt beyond reasonable doubt – discrepancies and inconsistencies in complainant’s evidence – appeal allowed – conviction quashed. Case 2: DEVELOPMENT APPEAL – dual occupancy – contentions resolved - covenant – variation of covenant – weight to be given to covenant - view sharing – community objections. Case 3: PROCEDURE – application for separate questions under UCPR 28.2 – where plaintiff seeks orders under Part 1C of the Civil Liability Act 2002 to set aside settlement agreements which may otherwise preclude the plaintiff from maintaining the balance of the proceedings – claimed benefit to defendants of plaintiff’s promise to “take no action”– statutory right to commence proceedings subsequently enacted – construction of s 7D(1) of the Civil Liability Act 2002 – overlap in issues and credit for hearing of separate questions and the balance of the proceedings – application for separate questions rejected.
Figure 2: Examples of Catchwords from different cases in NSW Caselaw.
A.2 Distribution of Citations
Figure 3 (Top) shows the frequency distribution of all $18,677$ citations in the data, and Figure 3 (Bottom) shows the top-20 most frequent citations.
Figure 3: (Top) Frequency distribution of unique cases in the data. The red vertical dashed line marks the last case with citation frequency of 9 or higher. (Bottom) Top-20 most frequently cited cases in the data.
A.3 Details of Instruction Fine-tuning
The instruction details and training configurations used for instruction fine-tuning are listed in Table 3 and Table 10. The re-rankers use the same training configurations.
Aspect Details
Instruction Predict the name of the case that needs to be cited in the text and explain why it should be cited.
Input test_set[i][’citation_text’].replace(test_set[i][’cited_case_name’], ’<CASENAME>’)
Output test_set[i][’citation_reason’] + <test_set[i][’citation’]>
Training Details
   GPU Single A100 GPU with 80G memory
   Optimizer adamw_torch
   Epochs 10
   Learning Rate 2e-4
   Quantization Using 8-bit quantization for LoRA training
LoRA Settings
   $r$ 16
   $\alpha$ (LoRA Alpha) 32
   Dropout 0.05
   Target Modules [up_proj, down_proj, gate_proj, k_proj, q_proj, v_proj, o_proj]
Table 3: Instruction fine-tuning and training details used for training our citation-specialised LLMs.
A.4 Prompt details for various experiments
All prompting details (or few-shot demonstrations when applicable) used in our experiments are provided in Table A.4-A.4.
Prompt template for the RoC Generation in our data
system_prompt = """ I have a text from a legal case document that includes a citation to another legal case. A case may be cited for different reasons. You will be provided with the text, the cited case name, and the cited case text. You must strictly follow these instructions: 1. If the text contains only case names and no further details or context related to the reason for citation, you must generate exactly: NOT ENOUGH INFORMATION. 2. If there is sufficient context in the text to show the reason why the case is cited in this text, you should summarize the reason with a detailed analysis. 3. You should put the reason in the first sentence, and the detailed analysis in the following sentences. 4. You should be conservative as much as possible, do not speculate the reason by yourself even if you may have some knowledge about the case. 5. Only provide the reason when the text contains enough information to get the reason. Otherwise, just generate NOT ENOUGH INFORMATION. """ prompt = """ Text: TEXT ------------------------ Cited Case Name: CASENAME ------------------------ Cited Case Text: CASETEXT ------------------------ Response: """
Table 4: The prompt template for the generation of Reason-of-Citation for cases in the data.
Prompt template used in the LLM-only experiments of Table 1
[…]
Table 5: The prompt template used for direct prompting of LLMs (general purpose and law specialised) in Table 1. Note, for our instruction-tuned LLMs we used the instructions and not these prompts.
[…]
system_prompt_few_shot = """ The following text description belongs to a case in the NSW Case Law. You will be given a brief text containing a cited case with a masked token <CASENAME>. Your task is to predict the citation reason of this case. Here are some examples for your reference: Text: This is not a case where it is appropriate for this Court to deal with the challenges to other findings of breaches of duty by MetLife in its consideration and determination of whether the TPD definition was satisfied. Many of those findings take account of his Honours conclusion in relation to MetLifes rejection of the lay witness material, and it would be difficult and artificial to deal with those finding on the hypothesis that this rejection involved no breach of duty: see <CASENAME> at [7] (Leeming JA, Basten and Gleeson JJA agreeing). Citation Reason: The cited case is referenced to support the argument that it would be inappropriate for the Court to address challenges to findings of breaches of duty without considering the context of those findings. Text: The third order sought was that the Tribunal "award exemplary costs against the Tribunal Registry." As the appellant correctly notes in his submissions the NCAT Act at s 60 sets out that the Tribunal may award costs in proceedings. The Tribunal has found on various occasions that non-parties can be the subject of costs orders: The Owners - Strata Plan No 79749 v Dunstan [2022] NSWCATAP 262; <CASENAME>. Citation Reason: The cited case is referenced to support the assertion that non-parties can be subject to costs orders in tribunal proceedings. Text: It does not seem to me that the appellant necessarily intends them in that manner, however that is the legal effect of the applications which are not properly brought before the Tribunal and cannot be determined in the appellants favour. It is the objectively ascertained purpose of the applications, and not the appellants subjective intent, which is relevant in that regard: <CASENAME> at [28]. Citation Reason: The cited case is referenced to emphasize the distinction between the subjective intent of the appellant and the objective legal effect of the applications. """ prompt = Text: + test_set[i][’citation_text’].replace(test_set[i][’cited_case_name’], ’<CASENAME>’)
Table 6: The prompt template used for generating auxiliary RoC (i.e., RoCaux) from LLMs in Table 1.
The prompt template used for LLMs in the RAG experiment of Table 1
[…]
Table 7: The prompt template used for re-ranking with LLMs in the RAG experiment—corresponding to Catch Words—in Table 1. Similar prompts are used for RoC Aggregations and Full Cases experiments.
Prompt template for merging the citation reasons of each case in re-ranker setting 2
system_prompt = """ You will be provided with a case from NSW Case Law, along with reasons it may be cited. Your task is to concisely summarize the reasons into a few sentences (2~3 sentences), if the original citation reasons are too long. """ prompt = """ Casename: CASENAME Citation Reasons: CITATIONREASON """
Table 8: The prompt template for merging the citation reasons of each case into text of similar length.
The prompt template used for generating the citation rationale in re-ranker setting 3
ranker_reason_sys_prompt = """ The following description belongs to a case in the NSW Case Law, you should predict the citation in the text. You will be given a brief text, 5 potential citations, their corresponding citation reasons and a ground-truth correct citation with its potential citation reasons. Note that the ground-truth citation may not be included among the five potential options. Your task is to predict the missing citation, provide a rationale before you draw a conclusion. Your predicted citation should be the ground-truth correct citation, but dont let on that you already know the answer (Dont mention anything about the ground-truth). The rationale should not be over 2~3 sentences. """ ranker_reason_prompt = """ Text: TEXT Potential Citations: CITATION1 Citation Reasons: CITATIONREASON1 CITATION2 Citation Reasons: CITATIONREASON2 CITATION3 Citation Reasons: CITATIONREASON3 CITATION4 Citation Reasons: CITATIONREASON4 CITATION5 Citation Reasons: CITATIONREASON5 Ground-truth Correct Citation: GROUNDTRUTH Ground-truth Citation Reasons: GTREASON """
Table 9: The prompt template used for generating the ground-truth citation rationale given candidate citations in re-ranker setting 3.
Instruction Predict the citation in the text.
Setting 1 Input
test_set[i][’citation_text’].replace(test_set[i][’cited_case_name’], ’<CASENAME>’) +
Potential Citations:
CITATION1
Citation Reasons: FIRST_CITATIONREASON1
CITATION2
Citation Reasons: FIRST_CITATIONREASON2
CITATION3
Citation Reasons: FIRST_CITATIONREASON3
CITATION4
Citation Reasons: FIRST_CITATIONREASON4
CITATION5
Citation Reasons: FIRST_CITATIONREASON5
Output <GOLD CITATION>
Setting 2 Input
[…]
Potential Citations:
CITATION1
Citation Reasons: MERGED_CITATIONREASON1
CITATION2
Citation Reasons: MERGED_CITATIONREASON2
CITATION3
Citation Reasons: MERGED_CITATIONREASON3
CITATION4
Citation Reasons: MERGED_CITATIONREASON4
CITATION5
Citation Reasons: MERGED_CITATIONREASON5
Output GOLD_CITATION_REASON + <GOLD CITATION>
Setting 3 Input
test_set[i][’citation_text’].replace(test_set[i][’cited_case_name’], ’<CASENAME>’) +
Potential Citations:
CITATION1
Citation Reasons: MERGED_CITATIONREASON1
CITATION2
Citation Reasons: MERGED_CITATIONREASON2
CITATION3
Citation Reasons: MERGED_CITATIONREASON3
CITATION4
Citation Reasons: MERGED_CITATIONREASON4
CITATION5
Citation Reasons: MERGED_CITATIONREASON5
Output RATIONAL + <GOLD CITATION>
Setting 4v1 Input
test_set[i][’citation_text’].replace(test_set[i][’cited_case_name’], ’<CASENAME>’) +
Potential Citations:
CITATION1
Citation Reasons: Top-1_CITATIONREASON1
CITATION2
Citation Reasons: Top-1_CITATIONREASON2
CITATION3
Citation Reasons: Top-1_CITATIONREASON3
CITATION4
Citation Reasons: Top-1_CITATIONREASON4
CITATION5
Citation Reasons: Top-1_CITATIONREASON5
Output RETRIEVED_TOP_1_CITATION_REASON + <GOLD CITATION>
Setting 4v2 Input
test_set[i][’citation_text’].replace(test_set[i][’cited_case_name’], ’<CASENAME>’) +
Potential Citations:
CITATION1
Citation Reasons: Top-1_CITATIONREASON1
CITATION2
Citation Reasons: Top-1_CITATIONREASON2
CITATION3
Citation Reasons: Top-1_CITATIONREASON3
CITATION4
Citation Reasons: Top-1_CITATIONREASON4
CITATION5
Citation Reasons: Top-1_CITATIONREASON5
Output GOLD_CITATION_REASON + <GOLD CITATION>
Table 10: The used instruction, input and output for the training of different re-rankers.